# A polynomial delay algorithm generating all potential maximal cliques in triconnected planar graphs

### v2

@@ Abstract
We develop a new characterization of potential maximal cliques of a triconnected planar graph and, using this characterization, give a polynomial delay algorithm generating all potential maximal cliques of a given triconnected planar graph. Combined with the dynamic programming algorithms due to Bouchitté and Todinca, this algorithm leads to a treewidth algorithm for general planar graphs that runs in time linear in the number of potential maximal cliques and polynomial in the number of vertices.
@@ 1 Introduction
Our treewidth algorithm for planar graphs is the first result where the planarity helps the exact treewidth computation in a non-trivial and essential manner.
@@
The simplicity of our characterization comes from the use of what we call a latching graph of a triconnected plane graph instead of radial graphs (for a definition, see [13] for example) or nooses that are used in conventional approaches for reasoning about separators in plane graphs. The latching graph $L_{G}$ of a biconnected plane graph $G$ is a multigraph obtained from $G$ by adding, in each face of $G$, every chord of the bounding cycle of the face. We observe that $L_{G}$ is simple if $G$ is triconnected (Proposition 2.5). Separators of $G$ correspond to cycles of $L_{G}$ in a similar way as they correspond to cycles of the radial graph of $G$. Unlike a radial graph, however, that is a bipartite graph on the set $V(G)$ of vertices and the set $F(G)$ of faces of the embedded graph $G$, a latching graph is a graph on $V(G)$ and simpler to reason with for some purposes. More specifically, we define a class of graphs we call steerings (Definition 3.1 in Section 3) and show that a vertex set $X$ of a triconnected plane graph $G$ is a PMC if and only if the subgraph of $L_{G}$ induced by $X$ is a steering. This characterization would be less simple if it was formulated in terms of radial graphs or nooses. It is not clear if the latching multigraph is useful for reasoning with plane graphs that are not triconnected. Adding this notion of latching graphs to the toolbox for triconnected planar graphs is one of our contributions.
@@ 2 Preliminaries
A vertex set $X$ of $G$ is a potential maximal clique (PMC for short) of $G$ if there is some minimal triangulation $H$ of $G$ such that $X$ is a maximal clique of $H$. We denote by $\Pi(G)$ the set of all PMCs of $G$. The following characterization of PMCs, which we heavily rely on, is due to Bouchitté and Todinca [8].
@@
A combinatorial representation of a plane graph $G$ is an indexed set $\{\pi_{v}\}_{v\in V(G)}$, where $\pi_{v}$ is a permutation on $N_{G}(v)$ such that $\pi_{v}(u)$ for $u\in N_{G}(v)$ is the neighbor of $v$ that comes immediately after $u$ in the clockwise order around $v$. Suppose we have two planar embeddings $G_{1}$ and $G_{2}$ of a graph $G$ and assume that $V(G_{1})=V(G_{2})=V(G)$ and moreover the isomorphisms between $G_{1}$ and $G$ as well as between $G_{2}$ and $G$ are identities. Let, for $i\in\{1,2\}$, the combinatorial representation of $G_{i}$ be $\{\pi_{i,v}\}_{v\in V(G)}$. We say that these two planar embeddings are combinatorially equivalent if either $\pi_{1,v}=\pi_{2,v}$ for every $v\in V(G)$ or $\pi_{1,v}=(\pi_{2,v})^{-1}$ for every $v\in V(G)$.
In this paper, we are mainly addressing triconnected planar graphs. It is known that, for a triconnected planar graph $G$, the planar embedding of $G$ is unique up to combinatorial equivalence [19]. For this reason, we will be dealing with triconnected plane graphs rather than triconnected planar graphs.
@@
Figure 1: (A) a triconnected plane graph (solid edges) and its latching graph (solid and broken edges)
(B) edge $\{u,v\}$ is drawn in more than one face if $u$ and $v$ separate the graph.
@@
Let $G$ be a triconnected plane graph and let $X\subseteq V(G)$ be such that $L_{G}[X]$ is a biconnected plane graph. Then, each $C\in{\mathcal{C}}_{G}(X)$ is contained in some region of $L_{G}[X]$. Moreover, each region of $L_{G}[X]$ contains at most one component in ${\mathcal{C}}_{G}(X)$.
Proof 2.10.
Since no edge of $G$ crosses any edge of $L_{G}$, each edge of $G[V(G)\setminus X]$ has its ends in the same region of $L_{G}[X]$. Therefore, for each $C\in{\mathcal{C}}_{G}(X)$, all the vertex of $C$ lie in the same region of $L_{G}[X]$. For the second statement, suppose a region $r$ of $L_{G}[X]$ contains more than one component in ${\mathcal{C}}_{G}(X)$. Then, there is some $x\in X$ on the boundary of $r$ that is adjacent to vertices from more than one component in ${\mathcal{C}}_{G}(X)$ contained in $r$. Let $e_{1}$, $e_{2}$, …, $e_{m}$ be the edges incident to $x$ lying in $r$, arranged in the clockwise ordering, and let $v_{i}$, $i\in\{1,\ldots,m\}$, be the other end of edge $e_{i}$. Since $r$ is a region of $L_{G}[X]$, none of $v_{i}$, $i\in\{1,\ldots,m\}$, belongs to $X$. Therefore, for some $i$, $1\leq i<m$, $v_{i}$ and $v_{i+1}$ belong to distinct components in ${\mathcal{C}}_{G}(X)$. But the face of $G$ to which $x$, $v_{i}$, and $v_{i+1}$ are incident is not incident to a vertex in $X$ other than $x$, since, if it did, the cycle bounding $r$ would have a chord in $L_{G}[X]$ that separates $r$, a contradiction. Therefore, there is a path of $G[V(G)\setminus X]$ between $v_{i}$ and $v_{i+1}$ around this face, also a contradiction. We conclude that $r$ contains at most one component in ${\mathcal{C}}_{G}[X]$.
@@
Proof 2.17.
Let $X$ be a PMC of $G$. We first show that $L_{G}[X]$ is biconnected. Consider an arbitrary vertex $x\in X$ in the PMC. We show that $L_{G}[X\setminus\{x\}]$ is connected. Let $y$ and $z$ be any two members of $X\setminus\{x\}$. Due to the third condition for PMCs, either $y$ is adjacent to $z$ in $G$ or there is some $C\in{\mathcal{C}}_{G}(X)$ such that $y,z\in N_{G}(C)$. In the former case, $y$ is adjacent to $z$ in $L_{G}[X]$. In the latter case, $S=N_{G}(C)$ is a minimal separator of $G$ and, therefore, $L_{G}[S]$ is a cycle. It follows that there is a subpath of the cycle $L_{G}[S]$ between $y$ and $z$ that does not contain $x$. Therefore, $L_{G}[X\setminus\{x\}]$ is connected for every $x\in X$ and hence $L_{G}[X]$ is biconnected.
To show that $L_{G}[X]$ is a plane graph, suppose that it is not the case. Due to Proposition 2.7, there is a face $f$ of $G$ such that $|V(f)\cap X|\geq 4$. Let $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$ be arbitrary four vertices in $V(f)$, listed in the order around $f$. If $x_{1}$ is not adjacent to $x_{3}$ in $G$, there is some $C\in{\mathcal{C}}_{G}(X)$ such that $x_{1},x_{3}\in N_{G}(C)$. In either case, there is a path $p_{1}$ of $G$ between $x_{1}$ and $x_{3}$ that intersects $X$ only at $x_{1}$ and $x_{3}$. Similarly, there is a path $p_{2}$ of $G$ between $x_{2}$ and $x_{4}$ that intersects $X$ only at $x_{2}$ and $x_{4}$. Since the Jordan curve obtained from $p_{1}$ by adding a curve in $f$ separates $x_{2}$ and $x_{4}$, $p_{1}$ and $p_{2}$ must intersect at a vertex of $G$. Therefore, there is some component $C\in{\mathcal{C}}_{G}(X)$ such that $x_{i}\in N_{G}(C)$ for every $i\in\{1,2,3,4\}$. But $S=N_{G}(C)$ cannot be a minimal separator, since $|S\cap V(f)|\geq 4$ and therefore $L_{G}[S]$ is not a cycle. This contradicts the assumption that $X$ is a PMC. Therefore, $L_{G}[X]$ is a plane graph.
Let $r$ be an arbitrary region of $L_{G}[X]$. If $r$ is non-empty and $C$ is the set of vertices lying in $r$, then $C\in{\mathcal{C}}_{G}(X)$ due to Proposition 2.9 and hence $S=N_{G}(C)$ is a minimal separator of $G$ since $X$ is a PMC of $G$. Due to Proposition 2.11, $L_{G}[S]$ is a cycle and bounds $r$. Suppose, on the other hand, that $r$ is empty. Then the cycle bounding $r$ has exactly three vertices because otherwise $G$ has a face $f$ such that $|V(f)\cap X|\geq 4$ and $L_{G}[X]$ would not be a plane graph due to Proposition 2.7. This cycle, having exactly three vertices, is trivially chordless.
@@ 3 Characterizing PMCs in triconnected planar graphs
Let $\gamma$ be a cycle. A subset $R$ of $V(\gamma)$ is a slot of $\gamma$ if $R$ is a singleton or an edge of $\gamma$. A socket of $\gamma$ is an unordered pair $\{R_{1},R_{2}\}$ of slots such that $R_{1}\cup R_{2}$ is not a slot. A vertex $s$ of $\gamma$ is a hinge of a socket $\{R_{1},R_{2}\}$ if $R_{1}\cup R_{2}$ is a subset of $N_{\gamma}(s)\cup\{s\}$.
A graph $H$ is a steering, if there is a bipartition $(S,P)$ of $V(H)$ such that $H[S]$ is a cycle, $N_{H}(P)$ is neither empty nor a slot of $H[S]$, and if $|P|\geq 2$ then the following conditions hold.
@@
    The unordered pair $\{N_{H}(u_{1})\cap S,N_{H}(u_{2})\cap S\}$ is a socket of $H[S]$.
@@
    $N_{H}(P\setminus\{u_{1},u_{2}\})\cap S$ is either empty or a singleton $\{s\}$ where $s$ is a hinge of the socket $\{N_{H}(u_{1})\cap S,N_{H}(u_{2})\cap S\}$.
We call $H$ an $(S,P)$-steering in this situation.
Remark 3.2.
It is important to note that the bipartition $(S,P)$ of $V(H)$ such that $H$ is an $(S,P)$-steering may not be unique in general. Also note that there is at most one hinge of a socket if $|S|\geq 5$ and at most two if $|S|=4$.
Figure 2 shows some steerings. Figure 3 shows two steerings with different $(S,P)$-partitions. For the second steering, the first bipartition gives rise to a hinge $s$ but the second bipartition does not involve a hing. Figure 4 shows some graphs that are not steerings.
Figure 2: Some steerings. White vertices belong to $P$, in a choice of the bipartition $(S,P)$ that works.
Figure 3: (A) A steering with three possible $(S,P)$-bipartitions. (B) A steering with two of the possible bipartitions. The first bipartition leads to a hinge $s$ while the second bipartition does not involve a hinge.
Figure 4: Some graphs that are not steerings. Neither the shown bipartition nor any other bipartition works.
@@
Every steering is planar.
Proof 3.4.
Let $H$ be a $(S,P)$-steering. A planar embedding of $H$ can be obtained in the following manner. First draw the cycle $H[S]$. Then choose one of the two faces of $H[S]$ and draw $H[P]$ arbitrarily in the interior of the chosen face. Let $\pi$ be the permutation on $S$ representing the clockwise order of $S$ around this chosen face. If $P$ is a singleton $\{u\}$, we can straightforwardly draw edges between $u$ and members of $S$ without crossings in accordance with the order prescribed by $\pi$.
Suppose $|P|\geq 2$ and let $u_{1}$ and $u_{2}$ be the ends of the path $H[P]$ and let $R_{i}=N_{H}(u_{i})\cap S$ for $i\in\{1,2\}$. Due to Condition 2 in the definition of steerings, $\{R_{1},R_{2}\}$ is a socket of $H[S]$. Due to Condition 3, $N_{H}(P\setminus\{u_{1},u_{2}\})\cap S$ is either empty or $\{s\}$ where $s$ is a hinge of the socket $\{R_{1},R_{2}\}$. Suppose the latter. Since $s$ is a hinge of the socket $\{R_{1},R_{2}\}$, $R_{1}$ is either $\{s_{1}\}$ or $\{s_{1},s\}$ and $R_{2}$ is either $\{s_{2}\}$ or $\{s_{2},s\}$, where $s_{1}$ and $s_{2}$ are two neighbors of $s$ on the cycle $H[S]$. We assume that $s_{1}$, $s$, and $s_{2}$ appear on $H[S]$ in this clockwise order around the region containing $P$. We order the edges between $P$ and $S$ in accordance with the clockwise order of $S$ prescribed by $\pi$: first the edges between $u_{1}$ and $S$, then the edges between $P\setminus\{u_{1},u_{2}\}$ and $s$, and finally the edges between $u_{2}$ and $S$. Following this order, we can draw edges between $P$ and $S$ without crossings. The drawing for the case $N_{H}(P\setminus\{u_{1},u_{2}\})\cap S=\emptyset$ is similar and simpler. We conclude that $H$ is planar.
@@
The first non-steering graph in Figure 4 does not satisfy Condition 1, having a face bounded by a cycle of length six going through the center vertex. This graph cannot be $L_{G}[X]$ for a PMC $X$ since the component of $G[V(G)\setminus X]$ that lies in the face in question would not be separated by a minimal separator. The second non-steering graph in the same figure has two vertices, the white vertex on the left and the right bottom black vertex, that are neither adjacent to each other nor share a common face. If $L_{G}[X]$ is this graph, then the third condition for $X$ to be a PMC in Lemma 2.1 would be violated. The third example has a similar problem.
@@
Proof 3.6.
Let $H=L_{G}[X]$ and let $(S,P)$ be a bipartition of $X$ such that $H$ is an $(S,P)$-steering. Since $L_{G}[S]=H[S]$ is a cycle, which we denote by $\gamma_{0}$, $L_{G}[S]$ has exactly two regions $r_{0}$ and $r$, one of which, say $r$ contains $P$. Since $r_{0}$ does not contain any vertex of $X$, $r_{0}$ is also a region of $L_{G}[X]$. Let $R=N_{L_{G}}(P)\cap S$ and let $E$ be the set of edges of $L_{G}[X]$ between $P$ and $S$. Due to the definition of a steering, $R$ is not a slot of $L_{G}[S]$ and hence $2\leq|R|\leq|E|$. The edges in $E$ divide the region $r$ of $L_{G}[S]$ into $k=|E|$ regions of $L_{G}[X]$ each bounded by a cycle consisting of two edges from $E$ together possibly with some edges of $L_{G}[S]$ and possibly with some edges of $L_{G}[P]$. Let $\Gamma$ denote the set of these $k$ cycles.
We show that each $\gamma\in\Gamma\cup\{\gamma_{0}\}$ is a chordless cycle of $L_{G}$. Since $S$ is a minimal separator of $G$, $\gamma_{0}=L_{G}[S]$ is a chordless cycle of $L_{G}$, due to Proposition 2.11. Let $\gamma$ be an arbitrary member of $\Gamma$. If $\gamma$ has a chord $\{u,v\}$ in $L_{G}$, then $\{u,v\}$ is an edge of $L_{G}[X]$ since $u,v\in X$. We show that no edge of $L_{G}[X]$ can be a chord of $\gamma$. We have three cases to consider: (1) $\{u,v\}$ is an edge of $\gamma_{0}$; (2) $\{u,v\}\in E$; (3) $\{u,v\}$ is an edge of the path $L_{G}[P]$. For (1), observe that the vertices in $V(\gamma)\cap S$ appear consecutively on $\gamma_{0}$. Therefore, for $\{u,v\}$ to be a chord of $\gamma$, $\gamma$ has to contain the subpath of $\gamma_{0}$ between $u$ and $v$ avoiding the edge $\{u,v\}$. However, this is impossible since $R$, being not a slot of $\gamma_{0}$, contains a vertex $w\not\in\{u,v\}$ and $E$ contains an edge incident to $w$. In cases (2) and (3), at most one of $u$ and $v$ belongs to $\gamma_{0}$ and therefore $\{u,v\}$ is not a chord of $\gamma_{0}$. For each $\gamma\in\Gamma$, if $u,v\in V(\gamma)$ then edge $\{u,v\}$ belongs to $\gamma$ and therefore is not a chord of $\gamma$. We conclude that every $\gamma\in\Gamma\cup\{\gamma_{0}\}$ is a chordless cycle of $L_{G}$.

### v3

@@ Abstract
We develop a new characterization of potential maximal cliques of a triconnected planar graph and, using this characterization, give a polynomial delay algorithm generating all potential maximal cliques of a given triconnected planar graph. Combined with the dynamic programming algorithm due to Bouchitté and Todinca, this algorithm leads to a treewidth algorithm for general planar graphs that runs in time linear in the number of potential maximal cliques and polynomial in the number of vertices.
@@ 1 Introduction
Our algorithm is the first one that explicitly exploits planarity for computing exact treewidth in a non-trivial manner.
@@
The simplicity of our characterization comes from the use of what we call a latching graph of a triconnected plane graph instead of the traditional tools for reasoning about separators in plane graphs: nooses or radial graphs, which are bipartite plane graphs representing the incidences of the vertices and faces (for a precise definition, see [16] for example). The latching graph $L_{G}$ of a biconnected plane graph $G$ is a multigraph obtained from $G$ by adding, in each face of $G$, every chord of the bounding cycle of the face. We observe that $L_{G}$ is simple if $G$ is triconnected (Proposition 2.6). Separators of $G$ correspond to cycles of $L_{G}$ in a similar way as they correspond to cycles of the radial graph of $G$. Unlike a radial graph, however, that is a bipartite graph on the vertices and the faces of the embedded graph $G$, a latching graph is a graph on $V(G)$ and simpler to reason with for some purposes. More specifically, we define a class of biconnected plane graphs we call steerings (Definition 3.1 in Section 3) and show that a vertex set $X$ of a triconnected plane graph $G$ is a PMC if and only if the subgraph of $L_{G}$ induced by $X$ is a steering. This characterization would be less simple if it was formulated in terms of radial graphs or nooses. It is not clear if the latching multigraph is useful for reasoning with plane graphs that are not triconnected. Adding this notion of latching graphs to the toolbox for triconnected planar graphs is one of our contributions.
@@ 2 Preliminaries
A vertex set $X$ of $G$ is a potential maximal clique (PMC for short) of $G$ if there is some minimal triangulation $H$ of $G$ such that $X$ is a maximal clique of $H$. We denote by $\Pi(G)$ the set of all PMCs of $G$. The following lemmas due to Bouchitté and Todinca [8] are essential in our reasoning about PMCs.
Lemma 2.1 (Bouchitté and Todinca [7]).
Let $X$ be a PMC of a graph $G$. Then, for every component $C$ of $G[V(G)\setminus X]$, $S=N_{G}[C]$ has a full component containing $X\setminus S$ and hence is a minimal separator. Moreover, for every minimal separator $S$ of $G$ such that $S\subseteq X$, we have $S\neq X$ and there is some component $C$ of $G[V(G)\setminus X]$ such that $S=N_{G}(C)$.
@@
A combinatorial representation of a plane graph $G$ is an indexed set $\{\pi_{v}\}_{v\in V(G)}$, where $\pi_{v}$ is a permutation on $N_{G}(v)$ such that $\pi_{v}(u)$ for $u\in N_{G}(v)$ is the neighbor of $v$ that comes immediately after $u$ in the clockwise order around $v$. Suppose we have two planar embeddings $G_{1}$ and $G_{2}$ of a graph $G$ and assume that $V(G_{1})=V(G_{2})=V(G)$ and moreover the isomorphisms between $G_{1}$ and $G$ as well as between $G_{2}$ and $G$ are identities. Let, for $i\in\{1,2\}$, the combinatorial representation of $G_{i}$ be $\{\pi_{i,v}\}_{v\in V(G)}$. We say that these two planar embeddings are combinatorially equivalent if either $\pi_{1,v}=\pi_{2,v}$ for every $v\in V(G)$ or $\pi_{1,v}=(\pi_{2,v})^{-1}$ for every $v\in V(G)$. In this paper, we mainly deal with triconnected planar graphs. It is known that, for a triconnected planar graph $G$, the planar embedding of $G$ is unique up to combinatorial equivalence [22]. For this reason, we will be dealing with triconnected plane graphs rather than triconnected planar graphs.
@@
Figure 1: (A) a triconnected plane graph (solid edges) and its latching graph (solid and broken edges). (B) edge $\{u,v\}$ is drawn in more than one face if $u$ and $v$ separate the graph.
@@
Let $G$ be a triconnected plane graph and let $X\subseteq V(G)$ be such that $L_{G}[X]$ is a biconnected plane graph. Then, each $C\in{\mathcal{C}}_{G}(X)$ is contained in some region of $L_{G}[X]$. Moreover, each region of $L_{G}[X]$ contains at most one component in ${\mathcal{C}}_{G}(X)$ and, if the region is incident to four or more vertices, exactly one component.
Proof 2.11.
Since no edge of $G$ crosses any edge of $L_{G}$, each edge of $G[V(G)\setminus X]$ has its ends in the same region of $L_{G}[X]$. Therefore, for each $C\in{\mathcal{C}}_{G}(X)$, all the vertices of $C$ lie in the same region of $L_{G}[X]$. For the second statement, suppose a region $r$ of $L_{G}[X]$ contains more than one component in ${\mathcal{C}}_{G}(X)$. Then, there is some $x\in X$ on the boundary of $r$ that is adjacent to vertices from more than one component in ${\mathcal{C}}_{G}(X)$ contained in $r$. Let $e_{1}$, $e_{2}$, …, $e_{m}$ be the edges incident to $x$ lying in $r$, arranged in the clockwise ordering, and let $v_{i}$, $i\in\{1,\ldots,m\}$, be the other end of edge $e_{i}$. Since $r$ is a region of $L_{G}[X]$, none of $v_{i}$, $i\in\{1,\ldots,m\}$, belongs to $X$. Therefore, for some $i$, $1\leq i<m$, $v_{i}$ and $v_{i+1}$ belong to distinct components in ${\mathcal{C}}_{G}(X)$. But the face of $G$ to which $x$, $v_{i}$, and $v_{i+1}$ are incident is not incident to a vertex in $X$ other than $x$, since, if it did, the cycle bounding $r$ would have a chord in $L_{G}[X]$ that separates $r$, a contradiction. Therefore, there is a path of $G[V(G)\setminus X]$ between $v_{i}$ and $v_{i+1}$ around this face, also a contradiction. We conclude that $r$ contains at most one component in ${\mathcal{C}}_{G}[X]$.
Finally, suppose a region $r$ of $L_{G}[X]$ is empty. Then $V(r)\subseteq V(f)$ for some face $f$ of $G$. We have $|V(r)|=3$, since otherwise $L_{G}[X]$ would not be a plane graph due to Proposition 2.8. Therefore, every region $r$ with $|V(r)|\geq 4$ is non-empty and hence contains exactly one component.
@@
Proof 2.18.
Let $X$ be a PMC of $G$. We first show that $L_{G}[X]$ is biconnected. Consider an arbitrary vertex $x\in X$ in the PMC. We show that $L_{G}[X\setminus\{x\}]$ is connected. Let $y$ and $z$ be any two members of $X\setminus\{x\}$. Due to the second condition for PMCs in Lemma 2.2, either $y$ is adjacent to $z$ in $G$ or there is some $C\in{\mathcal{C}}_{G}(X)$ such that $y,z\in N_{G}(C)$. In the former case, $y$ is adjacent to $z$ in $L_{G}[X]$. In the latter case, $S=N_{G}(C)$ is a minimal separator of $G$ and, therefore, $L_{G}[S]$ is a cycle due to Proposition 2.12. It follows that there is a subpath of the cycle $L_{G}[S]$ between $y$ and $z$ that does not contain $x$. Therefore, $L_{G}[X\setminus\{x\}]$ is connected for every $x\in X$ and hence $L_{G}[X]$ is biconnected.
To show that $L_{G}[X]$ is a plane graph, suppose that it is not the case. Due to Proposition 2.8, there is a face $f$ of $G$ such that $|V(f)\cap X|\geq 4$. Let $x_{1}$, $x_{2}$, $x_{3}$, and $x_{4}$ be arbitrary four vertices in $V(f)$, listed in the order around $f$. If $x_{1}$ is not adjacent to $x_{3}$ in $G$, there is some $C\in{\mathcal{C}}_{G}(X)$ such that $x_{1},x_{3}\in N_{G}(C)$. In either case, there is a path $p_{1}$ of $G$ between $x_{1}$ and $x_{3}$ that intersects $X$ only at $x_{1}$ and $x_{3}$. Similarly, there is a path $p_{2}$ of $G$ between $x_{2}$ and $x_{4}$ that intersects $X$ only at $x_{2}$ and $x_{4}$. Since the Jordan curve obtained from $p_{1}$ by adding a curve in $f$ separates $x_{2}$ and $x_{4}$, $p_{1}$ and $p_{2}$ must intersect at a vertex of $G$. Therefore, there is some component $C\in{\mathcal{C}}_{G}(X)$ such that $x_{i}\in N_{G}(C)$ for every $i\in\{1,2,3,4\}$. But $S=N_{G}(C)$ cannot be a minimal separator, since $|S\cap V(f)|\geq 4$ and therefore $L_{G}[S]$ is not a cycle. This contradicts the assumption that $X$ is a PMC, because of Lemma 2.1. Therefore, $L_{G}[X]$ is a plane graph.
Let $r$ be an arbitrary region of $L_{G}[X]$. If $r$ is non-empty and $C$ is the set of vertices lying in $r$, then $C\in{\mathcal{C}}_{G}(X)$ due to Proposition 2.10 and hence $S=N_{G}(C)$ is a minimal separator of $G$ due to Lemma 2.1 since $X$ is a PMC of $G$ . Due to Proposition 2.12, $L_{G}[S]$ is a cycle and bounds $r$. Suppose, on the other hand, that $r$ is empty. Then the cycle bounding $r$ has exactly three vertices because otherwise $G$ has a face $f$ such that $|V(f)\cap X|\geq 4$ and $L_{G}[X]$ would not be a plane graph due to Proposition 2.8. This cycle, having exactly three vertices, is trivially chordless.
@@ 3 Characterizing PMCs in triconnected planar graphs
Let $\gamma$ be a cycle. A subset $R$ of $V(\gamma)$ is a slot of $\gamma$ if $R$ is a singleton or an edge of $\gamma$. A graph $H$ is a steering, if there is a bipartition $(S,P)$ of $V(H)$ such that $H[S]$ is a cycle, $N_{H}(P)$ is neither empty nor a slot of $H[S]$, and if $|P|\geq 2$ then the following conditions hold.
@@
    No internal vertex of the path $H[P]$ is adjacent to any vertex in $S$.
@@
    For each end $t$ of the path, $N_{H}(t)\cap S$ is a slot of $H[S]$.
We call $H$ an $(S,P)$-steering in this situation. We call a steering a wheel if it is an $(S,P)$-steering for some bipartition $(S,P)$ of $V(H)$ such that $|P|=1$; otherwise it is a non-wheel steering.
Figure 2 shows some steerings. Figure 3 shows two steerings with different $(S,P)$-bipartitions. The example (B) is an $(S,P)$-steering for the second bipartition but not for the first bipartition. Figure 4 shows some graphs that are not steerings.
Figure 2: Some steerings. White vertices belong to $P$ and black vertices belong to $S$, in a choice of the bipartition $(S,P)$ that works. The first three are wheels while the remaining three are non-wheels.
Figure 3: (A) A steering with three possible $(S,P)$ bipartitions. Because of the first or the second bipartition, it is a wheel. (B) Also a wheel. For the first $(S,P)$ bipartition, it is not an $(S,P)$-steering since an internal vertex of the path on $P$ are adjacent to $s\in S$. The second $(S,P)$ bipartition shows that it is a wheel.
Figure 4: Some graphs that are not steerings. In the shown $(S,P)$ bipartition, the first example $H$ is not a steering since $N_{H}(P)\cap S$ is a slot. The second example is not a steering since $N_{H}(t)\cap S$ is not a slot for the right end $t$ of the path $H[P]$. The third example is not a steering since an internal vertex of the path $H[P]$ is adjacent to a vertex in $S$. It can be verified that these graphs are not $(S,P)$-steerings for any other $(S,P)$ bipartition. It is explained in the main text why any of these plane graphs cannot be $L_{G}[X]$ for a PMC $X$ of a triconnected plane graph $G$.
@@
Let $H$ be a steering. Then $H$ is biconnected and planar. Moreover, the planar embedding of $H$ is unique up to combinatorial equivalence.
Proof 3.3.
Let $H$ be a $(S,P)$-steering. We first confirm that $H$ is biconnected. Let $v$ be an arbitrary vertex of $H$. If $v\in S$, then $H[S\setminus\{v\}]$ is a path and $N_{H}(P)\cap(S\setminus\{v\})$ is non-empty as $N_{H}(P)\cap S$ is neither empty nor a singleton, so $H[V(H)\setminus\{v\}]$ is connected. If $v\in P$, then each component of $H[P\setminus\{v\}]$ has a vertex adjacent to a vertex in $S$ and hence $H[V(H)\setminus\{v\}]$ is connected.
We next show that $H$ is planar and has a unique planar embedding up to combinatorial equivalence. Suppose first that $|N_{H}(P)|=2$. Then the two members $s_{1}$ and $s_{2}$ of $N_{H}(P)$ are of degree three and all other vertices of $H$ are of degree two. We draw three paths between $s_{1}$ and $s_{2}$ without crossings and hence $H$ is planar. The combinatorial embedding of $H$ in this case essentially consists of the clockwise ordering $\pi_{1}$ of $N_{H}(s_{1})$ around $s_{1}$ and the clockwise ordering $\pi_{2}$ of $N_{H}(s_{2})$ around $s_{2}$. For each $i\in\{1,2\}$, we have two choices of $\pi_{1}$, one being the inverse of the other, and if we flip $\pi_{1}$ then $\pi_{2}$ is flipped. Therefore, there are only two possible combinatorial embeddings of $H$, which are equivalent to each other.
Suppose next that $|N_{H}(P)|\geq 3$. The planar drawing of $H$ can be obtained by first drawing the cycle $H[S]$, draw the path $H[P]$ in one of the faces of $H[S]$, and then drawing edges between vertices in $P$ and the vertices in $S$. If $|P|=1$ then these edges can certainly be drawn without crossings. Suppose $|P|\geq 2$ and let $t_{1}$ and $t_{2}$ be the two ends of the path $H[P]$. By the definition of an $(S,P)$-steering, $N_{H}[t_{i}]\cap S$ is a slot of the cycle $H[S]$ for $i\in\{1,2\}$. A crossing of edges would be unavoidable only if $N_{H}[t_{1}]\cap S$ and $N_{H}[t_{1}]\cap S$ are an identical edge of $H[S]$. But this is impossible since $N_{H}(P)$ would then be equal to this edge and hence would be a slot, violating the condition for $H$ to be an $(S,P)$-steering. In this case, where we are assuming that $|N_{H}(P)|\geq 3$, there is a vertex $v\in P$ of degree $\geq 3$ in $H$. If $|P|=1$ then $v$ is the unique member of $P$ and if $|P|\geq 2$ then $v$ is one of the ends $t_{1}$ and $t_{2}$ of the path $H[P]$. In either case $v$ is adjacent to at least two vertices in $S$. Since the ordering of $N_{H}(v)$ is constrained by the ordering of $S$ along the cycle $H[S]$, we have only two choices for the clockwise ordering $\pi_{v}$ of $N_{H}(v)$ around $v$, one being the inverse of the other. When we choose one of them, then the clockwise order of $S$ around the face containing $P$ is determined and this in turn determines the clockwise ordering $\pi_{u}$ of $N_{H}(u)$ around $u$ for every vertex $u$ of degree $\geq 3$ in $H$. When we flip the choice of $\pi_{v}$ then $\pi_{u}$ is flipped for every vertex $u$ of degree $\geq 3$ in $H$. Therefore, there are only two combinatorial embeddings of $H$, which are equivalent to each other.
From now on, when we refer to a steering $H$, we view $H$ as a plane graph rather than a combinatorial graph as originally defined.
@@
Also observe that if $L_{G}[X]$ is one of the non-steering graphs in Figure 4, then $X$ cannot be a PMC of $G$: in the first example, the component of $G[V(G)\setminus X]$ lying in the gray face would be a full component of $X$, violating the first condition of Lemma 2.2 for $X$ to be a PMC of $G$; in the second and third examples, the pair of checked vertices do not share a common face and therefore cannot belong to the neighborhood of a common component of $G[V(G)\setminus X]$, violating the second condition of Lemma 2.2 for $X$ to be a PMC of $G$.
@@
Proof 3.5.
Suppose $L_{G}[X]=H$ where $H$ is an $(S,P)$-steering. This equality means an equality as plane graphs, since we are viewing a steering as a biconnected plane graph, as we mentioned in the remark following Proposition 3.2.
Due to Proposition 2.10, each component in ${\mathcal{C}}_{G}(X)$ is contained in a region of $H$ and each region of $H$ contains at most one component in ${\mathcal{C}}_{G}(X)$.